What is the angle subtended by a vector A=i-j with x axis ?
Provide solution.
1
Expert's answer
2019-02-04T10:09:08-0500

A=i ̂-j ̂=A_x i ̂+A_y j ̂

So

A_x=1,A_y=-1

The angle between a vector and x axis

θ=tan^(-1)⁡〖A_y/A_x =tan^(-1)⁡〖(-1)/1=-45°〗 〗